6. On January 1, Wonder Water borrowed $300,000 for 5 years at 4.5% to finance expansion. Fixed principal payments are to be made quarterly beginning Mar 1. Below is an instalment schedule for Wonder Water. (13 marks)

WONDER WATER

INSTALMENT PAYMENT SCHEDULE- FIXED PRINCIPAL PAYMENTS

Interest Period

Cash Pmt

Interest Expense

Reduction of Principal

Principal

Jan 1

300,000

Mar 1

?

3,375

5,000

295,000

Jun 1

8,319

?

5,000

?

Sep 1

?

3,263

?

285,000

Dec 1

?

?

5,000

?

Instructions

(a) Determine the missing values (round to the nearest dollar).

(b) Prepare the journal entries for the payments made on March 1 and Sept 1.
Answer #1

Answer

Interest Period

Cash Pmt

Interest Expense

Reduction of Principal

Principal

Jan-01

$300,000

Mar-01

$8,375

$3,375

$5,000

$295,000

Jun-01

$8,319

$3,319

$5,000

$290,000

Sep-01

$8,263

$3,263

$5,000

$285,000

Dec-01

$8,206

$3,206

$5,000

$280,000

--Working

Cash Pmt

Interest Expense

Reduction of Principal

Principal

300000

=3375+5000

3375

5000

295000

8319

=8319-5000

5000

=295000-5000

=5000+3263

3263

5000

285000

=5000+3206

=285000*4.5%*0.25

5000

=285000-5000

  • Journal entries

Date

Accounts title

Debit

Credit

Mar-01

Notes Payable /Long term loans

$5,000

Interest Expense

$3,375

   Cash

$8,375

Sep-01

Notes Payable /Long term loans

$5,000

Interest Expense

$3,263

   Cash

$8,263